### v1

@@ VI Summary
For the first time, we validate a quantum machine learning model trained on real experimental data from the CMS detector, demonstrating that quantum approaches can extract meaningful physics insights in a real-world setting. These results establish the 1P1Q encoding as a viable and scalable data representation framework for quantum computing applications in particle physics. As quantum hardware continues to advance, the efficiency of the 1P1Q approach is tailored for more intricate and large-scale QML applications in high-energy physics, offering new perspectives for jet classification, anomaly detection, and beyond.
Acknowledgements

### v2

@@ Introduction.
For the first time, we validate a quantum machine learning model trained on real experimental data from the CMS detector, demonstrating that quantum approaches can extract meaningful physics insights in a real-world setting. These robust results establish the 1P1Q encoding as a viable and scalable data representation framework for quantum computing applications in particle physics. As quantum hardware continues to advance, the efficiency of the 1P1Q approach is tailored for more intricate and large-scale QML applications in high-energy physics, offering new perspectives for jet classification, anomaly detection, and measurements revealing the quantum mechanical nature of particle production.
Acknowledgements.
@@
Appendix
Training setup
The quantum circuits presented in this study are simulated and optimized using the Quantum Machine Learning library pennylane [35] with the lightning.gpu and lightning.kokkos devices. The trainable circuit parameters are optimized using the classical Adam Optimizer [36], with a scheduler that periodically decays the learning rate.
To train the QAE, we seek to maximize the fidelity between the output and input states in the subspace relevant for reconstruction:
$F$ $=\sum_{i=1}^{N_{\text{ref}}}\braket{T_{i}|R_{i}}$ $\forall~~~~T_{i}\in H^{T},R_{i}\in H^{\text{ref}}$ (7)
[…]
In the case of the VQC, we use the expectation value of the Pauli $Z$ observable on the target, which is the first qubit in the circuit. The final VQC prediction (Eq. 8) is arrived at by adding a classical, trainable bias term $b$ to the expectation value:
$f(x)$ $=\langle q(x)|Z|q(x)\rangle+b$ (8)
The circuit optimization is performed using the Mean Squared Error (MSE) between the prediction and the truth label as the loss function.
Datasets
[…]
To ensure the simulated jets closely resemble those reconstructed by the CMS detector, the detector effects are simulated with Delphes [41] using a simplified CMS detector configuration. Jets are clustered utilizing the anti-$k_{\mathrm{T}}$ algorithm [42] with a distance parameter $R=0.8$. As additional criteria, only jets with $p_{\mathrm{T}}\in[500,1000]\,$GeV and pseudorapidity $|\eta|<2$ are stored. For signal jets, additional quality requirements are imposed to ensure they fully contain the decay products of the initial particles.
[…]
For training the QAE, we use a total of 10,000 events for training and 2,500 events for validation. Inference is performed on 10,000 events of background and of each of the following class of events: $H\to b\overline{b}$, $H\to c\overline{c}$, $H\to gg$, $W\to q\overline{q^{\prime}}$, $Z\to q\overline{q}$ and $t\to bq\overline{q^{\prime}}$. The VQC is trained using 1,000 events and validated on 500 events, with an inference dataset of 10,000 events equally distributed between signal and background classes.
For training the QAE and VQC respectively, we first sample jets such that each class has a flat distribution in jet $p_{\mathrm{T}}$, in the range $[500,1000]$ GeV, so as to not bias the training towards the scale of the jet, allowing us to purely focus on jet substructure. The effect of this sampling can be seen in Figure 6. The jet samples used in the inference are also sampled to have the same flat distribution.
Figure 6: Effect of the $p_{\mathrm{T}}$ sampling on the input datasets used for training